OpenStudy (anonymous):

The probability of a spinner stopping on 3 is . The probability of the spinner stopping on 6 is . What is the probability of the spinner stopping on 3 or 6? A. B. C. D.

OpenStudy (anonymous):

@amorfide

OpenStudy (anonymous):

medal and fan

Miracrown (miracrown):

What do the ABCD = ?

OpenStudy (anonymous):

I just need the nswer

Miracrown (miracrown):

And also I cannot see the numbers after 3 and 6 ...

Miracrown (miracrown):

I am not here to provide you with answers. I'll guide you through the steps, so then you can come up with the answer.

OpenStudy (anonymous):

1/8 is the number srry

OpenStudy (amorfide):

if you are saying that the probability to spin a 3 is 1/8 and the probability to spin a 6 is 1/8 the probability to spin a 3 or a 6 is the added probabilities so 1/8 + 1/8

OpenStudy (anonymous):

a is 2/8 b is 1/16 c is 1/8 and D is 1/64

OpenStudy (anonymous):



Miracrown (miracrown):

We can just add these two probabilities: 1/8+1/8=1/4. That is, the probability that the spinner stops at 3 or 6 is 1/4.

OpenStudy (amorfide):

add up the probabilities

OpenStudy (anonymous):

oh I just need the answer medal and fan

Miracrown (miracrown):

What about a no? I just showed you the steps and now what you have to do is just simply add up the probability and then you'll arrive to the answer.

OpenStudy (amorfide):

\[\frac{ a }{ b } + \frac{ c }{ b }=\frac{ a+c }{ b }\] where a=1 c=1 b=8

OpenStudy (anonymous):

so it is 1/8